OpenStudy (anonymous):

we are having trouble with a domain question, how would we find the domain of y = the square root of tan x

OpenStudy (anonymous):

i needdddd hellppppp p p p

OpenStudy (anonymous):

how

OpenStudy (anonymous):

The domian is all values of which tan x is positive

OpenStudy (anonymous):

remember we cannot have negatives under the square root

OpenStudy (anonymous):

so how would it be - infin, to infin

OpenStudy (anonymous):

no it wouldnt

OpenStudy (anonymous):

0 to infinity

OpenStudy (anonymous):

it would be:x=[kpi,kpi+pi/2), \[k \epsilon z\]

OpenStudy (anonymous):

what is k

OpenStudy (anonymous):

xan u explant howyou got that answer

OpenStudy (anonymous):

what values of tan x give us positive values??, where is it negative? use unit circle

OpenStudy (anonymous):

Since we are using the unit circle to determine where tan is positive and negative, we have a huge domain , k can represent any number, 0,1,2,3,4,5......

OpenStudy (anonymous):

look at this: \[\sqrt{tanx}>0\]

OpenStudy (anonymous):

x has to be in the 1sr or 3rd quadrant, since tan is positive in those quadrant

OpenStudy (anonymous):

at pi over 2

OpenStudy (anonymous):

no , at pi/2 tanx is undefiened

OpenStudy (anonymous):

3pi 0ver 2

OpenStudy (anonymous):

\[0+k \pi \le x <\pi/2 +k \pi\] for all integers k

OpenStudy (anonymous):

satisfied?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

satisfied?

OpenStudy (anonymous):

want t still help or no

OpenStudy (anonymous):

yeah, what else you got

OpenStudy (anonymous):

solve the following by factoring and making appropriate sign charts.\[x ^{2}-16>0\]

OpenStudy (anonymous):

this is a difference of squares

OpenStudy (anonymous):

basically all i need is a sign chart buddy boy

OpenStudy (anonymous):

factor it, then see what your signs will be.

OpenStudy (anonymous):

then you first have to factor and get the values at which that polynomial is zero buddy boy

OpenStudy (anonymous):

im a girllllllllllllllllllllllll genius

OpenStudy (anonymous):

okay, how ma i suppose to know that girly

OpenStudy (anonymous):

did you factor yet

OpenStudy (anonymous):

Once you factor you should get three intervals, that will help you form your preciouse chart

OpenStudy (anonymous):

yes x greater than 4

OpenStudy (anonymous):

hold on this is a differnece of square you should get (x+4)(x-4)

OpenStudy (anonymous):

ok i see now sorry boiut that! so x equals 4 and negaative 4

OpenStudy (anonymous):

right, now you must form three intervals

OpenStudy (anonymous):

and how do you do that? i have never done sign charts before

OpenStudy (anonymous):

look i did it for you girly

OpenStudy (anonymous):

thank you im trying to zoom on it

OpenStudy (anonymous):

i can see it perfectly

OpenStudy (anonymous):

yeah me too, so what do we do with the chart?? plug in random numbers? can u do it exactly how you would do it on paper please?

OpenStudy (anonymous):

sure i can spend my time and do you problem exactly as it would look on paper

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

here you go buddy

OpenStudy (anonymous):

awe thank you bud! do you know how to do piecewise functions too?

OpenStudy (anonymous):

ill give it a shot

OpenStudy (anonymous):

whats the problem?